### v2

# CheXalign: Preference fine-tuning in chest X-ray interpretation models without human feedback

Abstract
Radiologists play a crucial role in translating medical images into actionable reports. However, the field faces staffing shortages and increasing workloads. While automated approaches using vision-language models (VLMs) show promise as assistants, they require exceptionally high accuracy. Most current VLMs in radiology rely solely on supervised fine-tuning. Meanwhile, additional preference fine-tuning in the post-training pipeline has become standard practice in the general domain. The challenge in radiology lies in the prohibitive cost of obtaining radiologist feedback at scale. To address this challenge, we propose an automated pipeline for preference feedback, focusing on chest X-ray radiology report generation (RRG). Specifically, our method leverages publicly available datasets containing pairs of images and radiologist-written reference reports with reference-based metrics, or Judges, eliminating the need for additional radiologist feedback. We investigate reward overoptimization via length exploitation in this setting and introduce a length-controlled version of the GREEN score. Our best-performing setup achieves state-of-the-art CheXbert scores on the MIMIC-CXR dataset for the RRG task while on average maintaining robust performance across six additional image perception and reasoning tasks.
CheXalign: Preference fine-tuning in chest X-ray interpretation models without human feedback
Dennis Hein heind@stanford.edu                        Zhihong Chen zhihongc@stanford.edu                        Sophie Ostmeier sostm@stanford.edu
Justin Xu xujustin@stanford.edu                        Maya Varma mvarma2@stanford.edu                        Eduardo Pontes Reis edreis@stanford.edu
Arne Edward Michalson arne64@stanford.edu                        Christian Bluethgen bluethgen@stanford.edu                        Hyun Joo Shin lamer-22@yuhs.ac
Curtis Langlotz langlotz@stanford.edu                        Akshay S Chaudhari akshaysc@stanford.edu
1 Introduction
X-rays are one of the most frequently collected imaging studies in clinical practice, with the advantages of wide availability, cost-effectiveness, and low radiation dose. Chest X-rays (CXR) are used for diverse purposes, with approximately 1.4 billion diagnostic X-ray examinations collected per year in the world (PAHO, 2012; Organization et al., 2016; Cid et al., 2024). The amount and significance of CXRs can pose a burden for radiologists and a potential negative impact for patients without timely interpretation, especially for those containing critical lesions (Ruutiainen et al., 2013; Hanna et al., 2017; Bruls and Kwee, 2020; Bhargavan et al., 2002; Lyon et al., 2015; Rimmer, 2017).
Recent strides in generative vision-language models (VLMs) hold promising implications for this high-stakes and low-data field (Liu et al., 2024a; Radford et al., 2021). Typically pre-trained using image-text contrastive learning and supervised fine-tuned using causal language modeling (a.k.a. next-token prediction), recent VLMs have started to demonstrate promising performance in CXR interpretation and radiology report generation (RRG) (Chen et al., 2024a; Bannur et al., 2024). In high-stakes fields like radiology, where accurate medical descriptions directly influence disease diagnosis and treatment decisions, the generated outputs must maintain high factual accuracy to ensure patient safety.
However, recent studies have shown that supervised fine-tuning (SFT) might be insufficient in the post-training process. For example, Hong et al. (2024) illustrate the limitation of SFT by training on a preference dataset, containing “good” and “bad” completions. By tracking the log probabilities of each during the course of training, they show that the log probabilities of the bad completions inadvertently increase alongside the good completions. Preference fine-tuning methods, such as reinforcement learning from human feedback (RLHF) (Ziegler et al., 2020; Stiennon et al., 2020; Ouyang et al., 2022), using Proximal Policy Optimization (PPO) (Schulman et al., 2017) or REINFORCE (Williams, 1992), and direct alignment algorithms (DAAs), such as Direct Preference Optimization (DPO) (Rafailov et al., 2023), effectively alleviate this problem by employing a negative gradient to lower probabilities of “bad” completions (Tajwar et al., 2024). In fact, most recent large language models (LLMs) (Ouyang et al., 2022; Bai et al., 2022a; Touvron et al., 2023; Jiang et al., 2024; Team et al., 2024) include some form of preference fine-tuning in their post-training pipeline. Yet, this approach has not yet been investigated within the medical vision-language domain.
Figure 1: Overview of the CheXalign pipeline. For a given dataset containing CXRs and radiologist written references (e.g., MIMIC-CXR), we obtain preference pairs using some automated reference-based metric (e.g., GREEN or the BERTScore), and then optimize the policy using light-weight DAAs (e.g., DPO or KTO).
The primary challenge hindering the application of preference fine-tuning in the post-training of VLMs in radiology is the prohibitive cost of obtaining radiologist preferences at scale. To overcome this obstacle, we introduce CheXalign, an automated pipeline for generating preference data for the crucial RRG task. Specifically, we leverage the availability of reference reports written by radiologists in a clinical setting within large, publicly available, datasets such as MIMIC-CXR (Johnson et al., 2019) and CheXpert Plus (Chambon et al., 2024). This allows us to use reference-based, uni-modal, metrics, such as GREEN (Ostmeier et al., 2024), a recent state-of-the-art LLM-based metric for evaluating CXR reports, to annotate generated reports in a factually grounded fashion. An overview of our preference fine-tuning pipeline is available in Fig. 1. Our approach enables us to obtain high-quality preference datasets in a fully automated and scalable manner. Using our proposed method, we systematically study how DAAs can be used to enhance the clinical efficacy of medical VLMs without any additional radiologist feedback. Our contributions are as follows:
  1. 1.
    We introduce an automated pipeline for preference pair generation in RRG models, circumventing the prohibitively expensive task of obtaining preference feedback from radiologists at scale.
  2. 2.
    We systematically evaluate and benchmark the proposed pipeline using different reference-based metrics, DAAs, and RRG models. Our findings indicate that the RRG performance can be improved even when using inexpensive, general domain, natural language generation (NLG) metrics for preference pair generation.
  3. 3.
    Using our proposed pipeline, we obtain new start-of-the-art CheXbert scores on the MIMIC-CXR data for the RRG task.
  4. 4.
    We study reward overoptimziation via length exploitation, and introduce the length-controlled GREEN score.
  5. 5.
    We benchmark our models post alignment on set of diverse additional image perception and reasoning tasks to assess whether there is an alignment tax in this setting.
2 Related Works
VLMs (Radford et al., 2021; Li et al., 2021, 2022, 2023a; Liu et al., 2024a) are a multi-modal extension to LLMs. In this setting, the prompt $x$ contains images and/or text. Typical tasks include Vision Question Answering (VQA) and image captioning (e.g., RRG in the field of radiology). There are also a line of works to extend VLMs to the medical domain (Thawkar et al., 2023; Hyland et al., 2023; Chaves et al., 2024; Tu et al., 2024; Bannur et al., 2024; Chen et al., 2024a, b; Lee et al., 2024b; Jin et al., 2024) which mainly focus on CXR interpretation and RRG due to the wide availability of public datasets (Johnson et al., 2019; Chambon et al., 2024). However, even with strong LLMs and vision-backbones, VLMs have been observed to “hallucinate” and produce outputs that are not factually grounded in the image (Zhou et al., 2024). Such hallucinations represent a significant risk in high-stakes healthcare fields such as radiology. Similar to Zhou et al. (2024), we pose the problem of hallucinations as an alignment problem and propose tackling it via preference fine-tuning. As noted above, preference fine-tuning, although now a standard part of the post-training pipeline for general domain LLMs, remains largely unexplored in the context of medical VLMs.
RLHF (Ziegler et al., 2020; Stiennon et al., 2020; Ouyang et al., 2022) is a very effective technique for aligning LLMs and VLMs with human preferences. It was instrumental in the development of frontier models such as ChatGPT and Bard (Lee et al., 2024a). RLHF involves two stages: reward modeling, where a reward model is learned on the preference data, and alignment, where reinforcement learning (RL) algorithms are used to optimized the proxy reward. While extremely effective, RLHF is very computationally heavy and can be finicky for non-experts. Relatively recently, a new class of algorithms called DAAs (Rafailov et al., 2024) have become increasingly popular. [^1] Used more loosely than in  Rafailov et al. (2024). This class of algorithms re-parameterize the reward model via a change-of-variables using the closed-formed solution to the RLHF objective, effectively bypassing both the reward modeling and RL stages. This yields alignment algorithms that remain performant yet computationally more lightweight and significantly easier to implement. DPO (Rafailov et al., 2023) was the first in this category and remains one of the most popular versions. After the advent of DPO, a large variety of DAAs have been suggested (Azar et al., 2023; Park et al., 2024; Ethayarajh et al., 2024; Hong et al., 2024). A brief introduction to RLHF and DAAs is available in §A. Another dimension for improved efficiency is reinforcement learning from AI feedback (RLAIF), first introduced in Bai et al. (2022b). Replacing humans with LLMs leads to significant reductions in cost, making it much more scalable, while maintaining high quality Lee et al. (2024a). In this work, we will leverage these advancements for scalable, fully automated, preference data generation and computationally lightweight alignment.
Since the reward model in the RLHF objective is learned, it is an imperfect proxy of the ground truth reward. As this proxy is optimized, ground truth performance might saturate or even deteriorate. [^2] As per Goodhart’s law: “When a measure becomes a target, it ceases to be a good measure.” (Gao et al., 2023). This reward overoptimization, or hacking, phenomena was first studied in Gao et al. (2023) for RLHF. Despite not fitting an explicit reward model, similar behavior has been observed empirically for DAAs (Rafailov et al., 2024). In particular, length exploitation, the tendency to learn to produce excessively verbose completions, is one common dimension of reward overoptimization, observed in both RLHF and for DAAs. For instance, Park et al. (2024) showed that DPO amplifies minor verbosity bias embedded in the preference data. In this work, we explore this phenomenon in the context of preference fine-tuning of RRG models.
3 Methodology
3.1 RRG Preference Fine-tuning without Human Feedback
Expert human feedback from radiologists is the gold standard for preference data generation and evaluation for the RRG task. However, scaling is impractical, if not unfeasible, due to the limited availability of radiologists for large-scale annotation tasks. In the general domain, it is common to leverage LLMs for cost effective preference data generation (Bai et al., 2022b; Dubois et al., 2023; Lee et al., 2024a)Zheng et al. (2023) categorized “LLM-as-a-Judge” evaluation methods into pairwise, single answer, and reference-guided grading. Pairwise grading being the most common in the general domain both for preference data generation (Dubois et al., 2023; Lee et al., 2024a) and evaluation (Zheng et al., 2023; Dubois et al., 2024).
These existing methods, however, are tailored for uni-modal, general-domain LLMs and do not directly apply to our multi-modal setting, which involves both visual and textual data. Moreover, factual grounding is essential in RRG to ensure clinical reliability. To overcome these challenges, we propose using reference-based grading, leveraging publicly available datasets that contain paired prompts—including images—and radiologist-written reference reports. This abundance of high-quality references allows us to provide factually grounded annotations without the need for a multi-modal metrics, or “Judge”, setting our approach apart from prior studies of preference alignment of VLMs, such as Sun et al. (2024). A comparison of reference-free and reference-based metrics, or “Judges”, in this setting is available in Fig. 2. For a given Judge, we obtain preference pairs by repeat sampling from the SFT baseline. Canonical alignment algorithms, such as DPO, can then be used to preference fine-tune the model.
Figure 2: Illustration of preference pairs generation using a reference-free multi-modal Judge (left) and a reference-based uni-modal Judge (right). Human (radiologist) feedback is multi-modal in nature but prohibitively expensive at scale. On the other hand, using LLM, or metrics-based, Judges is highly scalable. However, a high-quality multi-modal Judge is difficult to obtain. For the RRG task, we propose leveraging large publicly available datasets, containing CXRs and radiologist written reference reports, as this enables scalable, factually grounded, preference data generation using reference-based metrics.
3.2 Evaluation
Since we have radiologist-written reference available, it is possible to employ standard general domain NLG metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and the BERTScore (Zhang et al., 2019). However, these NLG metrics may not be able to differentiate between subtle nuances that are clinically relevant. Thus, we include two validated, radiology-specific, and clinically relevant, metrics: GREEN (Ostmeier et al., 2024) and CheXbert scores (Smit et al., 2020). GREEN is a state-of-the-art metric for radiology report evaluation, based on a single-answer reference-guided LLM-as-a-Judge mechanism. CheXbert scores are a clinical efficacy metrics based on extracting 14 labels [^3] Enlarged Cardiomediastinum, Cardiomegaly, Lung Opacity, Lung Lesion, Edema, Consolidation, Pneumonia, Atelectasis, Pleural Effusion, Pneumothorax, Pleural Other, Fracture, Support Devices, No Finding. using the CheXbert labeler (Smit et al., 2020) from the candidate and reference reports.
Initial experimentation illustrated that reward hacking via length exploitation might be a concern in our setting when using the GREEN score as Judge. To counteract this, we propose a simple heuristic approach to explicitly control for length:
[…]
where rel_verbosity denotes the relative verbosity (length in words) of the candidate report compared to the reference report. We call this metric length-controlled GREEN (LC-GREEN).
4 Experimental Details
4.1 SFT Baselines
We adopt CheXagent (Chen et al., 2024a) (8B) as a representative example of a state-of-the-art, open source, English language, VLM for the RRG task. It has been trained in a canonical way by first adapting an LLM, Mistral-7B (Jiang et al., 2023), to medical text by continued pre-training. Second, a vision encoder, EVA-CLIP-g (Sun et al., 2023),was adapted via vision pre-training, using contrastive learning on CXR image-text pairs. Third, the two modalities were merged by training a vision-language bridger, or adapter network, keeping the LLM and vision encoder frozen. Finally, the model was instruction-tuned for a range of tasks, including RRG. Concurrent to this work, CheXagent has been greatly improved upon with CheXagent-2 (Chen et al., 2024b) (3B). CheXagent-2 adopted a fine-tuned SigLIP Zhai et al. (2023) model [^4] https://huggingface.co/StanfordAIMI/XraySigLIP__vit-l-16-siglip-384__webli. as the vision encoder and a fine-tuned Phi-2 Li et al. (2023b) model [^5] https://huggingface.co/StanfordAIMI/RadPhi-2. (2.7B) as the language decoder. For the image-text connector, instead of using the attention layer as in CheXagent, it uses LLaVA-style MLP connector Liu et al. (2024b). CheXagent-2 additionally use indications, which provide clinical context to the patient being imaged, to aid in the RRG task.
4.2 Datasets
We use the MIMIC-CXR (Johnson et al., 2019) dataset for training, validation and testing. The image-report pairs consists of one or two CXRs and the corresponding free-text findings section. CheXagent-2 additionally use the indications. For CheXagent, we randomly sample 80k examples as our training data. For CheXagent-2, we opted to used the full MIMIC-CXR training set (148k examples). To test robustness for the RRG task, we additionally include test data from the CheXpert Plus (Chambon et al., 2024) dataset. Moreover, to evaluate whether there is an alignment tax, we additionally evaluate our aligned models on six additional CXR tasks: view classification, binary image classification, single disease identification, multi disease identification, VQA, and image-text reasoning, using test data from five additional datasets RSNA (Shih et al., 2019), SIIM (American College of Radiology, 2019), OpenI (Demner-Fushman et al., 2016), SLAKE (Liu et al., 2021), and Rad-Restruct (Pellegrini et al., 2023) datasets. All datasets are in English.
4.3 Preference Data
Algorithm Objective Preference Reference Length Relative
pairs controlled wall-clock time
DPO $-\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{c}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{c}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{r}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{r}\mid x\right)}\right)$checkmark $\checkmark$ $\times$ 1.0
LC-DPO $-\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{c}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{c}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{r}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{r}\mid x\right)}+\alpha(|y_{c}|-|y_{r}|)\right),\alpha>0$checkmark $\checkmark$ $\checkmark$ 1.0
IPO $\left(\log\frac{\pi_{\theta}(y_{c}|x)}{\pi_{\text{ref}}(y_{c}|x)}-\log\frac{\pi_{\theta}(y_{r}|x)}{\pi_{\text{ref}}(y_{r}|x)}-\frac{1}{2\tau}\right)^{2}$checkmark $\checkmark$ $\times$ 1.0
KTO $-\lambda_{c}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{c}|x)}{\pi_{\text{ref}}(y_{c}|x)}-z_{\text{ref}}\right)+\lambda_{r}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{r}|x)}{\pi_{\text{ref}}(y_{r}|x)}\right),$$\times$ $\checkmark$ $\times$ 2.2
$\text{where}\,\,z_{\text{ref}}=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\beta\mathbb{D}_{\text{KL}}\left(\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right)]$
ORPO $-\log p_{\theta}(y_{c}|x)-\lambda\log\sigma\left(\log\frac{p_{\theta}(y_{c}|x)}{1-p_{\theta}(y_{c}|x)}-\log\frac{p_{\theta}(y_{r}|x)}{1-p_{\theta}(y_{r}|x)}\right),$checkmark $\times$ $\times$ 0.7
$\text{where}\,\,p_{\theta}(y|x)=\exp\left(\frac{1}{|y|}\log\pi_{\theta}(y|x)\right)$
Table 1: Overview of the DAAs considered in this paper. Preference pairs indicates whether the method requires paired data of chosen/rejected or only binary feedback indicating whether a completion is desirable/undesirable. Reference indicates whether an additional reference model is loaded during training. Length controlled indicates whether the objective directly controls for the length of the completions in order to mitigate reward hacking via length exploitation. Relative wall-clock time, is measured as the total time to train one epoch, relative to DPO.
We evaluate two reference-based Judges: GREEN (Ostmeier et al., 2024) and the BERTScore (Zhang et al., 2019). We obtain our preference data as follows: 1) for each example in the training data, we prompt the SFT baseline $N=4$ times; 2) we get the score for each of the generated reports, compared with the corresponding singular reference; 3) we set the chosen and rejected completions as the highest and lowest scores, omitting the observation if all $N=4$ scores are equivalent. For CheXagent, this rejection rule results in the rejection of 1,246 (1.6%) examples for GREEN and 31 (0.04%) examples for the BERTScore. This is similar for CheXagent-2, with 3744 (2.53%) examples and 147 (0.10%) examples rejected for GREEN and the BERTScore, respectively. Summary statistics for the chosen and rejected subsets are available in Table 8. We also report summary statistics of the length (in words) of the generated reports. Notably, the difference, or spread, in average length of the chosen and rejected subsets is slightly more pronounced for GREEN than for the BERTScore, 6.9% compared to 5.8% and much more significant overall for CheXagent-2 with a difference of 19.3% and 17.1%, respectively.
Figure 3: Average length against average GREEN and BERTScore for all aligned policies on the MIMIC-CXR validation set using GREEN (left) and the BERTScore (right) as Judges.
4.4 Alignment Algorithms
We opt for a representative subset of available, offline, DAAs. DPO is the original DAA and serves as our baseline. In addition to DPO, we consider: 1) Length-controlled DPO (LC-DPO) (Park et al., 2024), as an example of a DAA with explicit length regularization. LC-DPO is an elegant extension of DPO, with an additional hyperparameter $\alpha$ which controls the strength of an additional length regularization term. Setting $\alpha=0$ reverts the objective to that in DPO. 2) Identity Preference Optimization (IPO) (Azar et al., 2023) as an example of a DAA with generalized preference, relaxing the assumption of the Bradley-Terry model used in DPO. The authors argue that this helps mitigate over-fitting issues even when preferences are transitive. Relatively recent work has shown that IPO indeed seems to be less prone to reward overoptimization (Rafailov et al., 2024). 3) Kahneman-Tversky optimization (KTO) (Ethayarajh et al., 2024) as an example of a DAA that does not require preference pairs, but instead only binary feedback on whether a completion is desirable or undesirable. This type of data is much more ubiquitous in practice. In addition, for any given dataset of preference pairs, KTO provides twice the number of examples. 4) Odds-Ratio Preference Optimization (ORPO) (Hong et al., 2024), almost outside of the definition of DAAs, is not based on the RLHF objective and instead appends an additional penalty directly to the negative log likelihood used in SFT. This adds a “negative gradient”, using the terminology in Tajwar et al. (2024), which will help reduce the log probabilities of rejected completions.
An overview of all DAAs considered in this paper is available in Table 1. More implementation details are available in §B.2.666Due to compute constraints, we only include results for the baseline alignment algorithm, DPO, for CheXagent-2.
5 Results and Analyses
5.1 Length Exploitation
We investigate the issue of length exploitation by plotting the average GREEN and BERTScore against average lengths of the resulting radiology reports using GREEN and the BERTScore, respectively, as Judges for preference data generation. Results for single runs on the MIMIC-CXR validation set, including all configurations considered, are available in Fig. 3. Save for a IPO run in the lower right corner, there seems to be a positive correlation between average GREEN and average length for CheXagent. In fact, for all DAAs, except for ORPO, there is an indication for a trade-off between added verbosity and GREEN. In addition, such a trade-off is not observed for the BERTScore, shown to the right in Fig. 3. Except for a very tight clustering around 0.87-0.88, there are no clear trends.
GREEN LC-GREEN
Method Avg. length Rel. verbosity Avg. length Rel. verbosity
CheXagent 55.8 55.8
+DPO 140.2 2.51 68.7 1.23
+LC-DPO 111.5 2.00 53.5 0.96
+IPO 100.2 1.80 58.0 1.04
+KTO 71.3 1.28 55.9 1.00
+ORPO 63.1 1.13 63.1 1.13
CheXagent-2 46.8 46.8
+DPO 60.9 1.30 54.7 1.17
Reference 58.4 1.05 58.4 1.05
Table 2: Average length of reports on the MIMIC-CXR validation set of best performing configurations according to GREEN and GREEN-LC, respectively. Relative verbosity is relative to the SFT baseline.
MIMIC-CXR CheXpert Plus
Model Judge GREEN ()($\uparrow$)( ↑ ) LC-GREEN ()($\uparrow$)( ↑ ) BERTScore()($\uparrow$)( ↑ ) GREEN ()($\uparrow$)( ↑ ) LC-GREEN ()($\uparrow$)( ↑ ) BERTScore()($\uparrow$)( ↑ )
CheXagent 0.249 0.218 0.856 0.248 0.202 0.851
+DPO 0.325 (30.6) 0.263 (20.5) 0.862 (0.65) 0.322 (29.9) 0.222 (9.69) 0.859 (0.91)
+LC-DPO 0.320 (28.7) 0.288 (32.0) 0.864 (0.90) 0.330 (32.8) 0.268 (32.3) 0.861 (1.26)
+IPO GREEN 0.326 (30.9) 0.282 (29.3) 0.863 (0.84) 0.334 (34.6) 0.267 (31.9) 0.861 (1.15)
+KTO 0.328 (31.9) 0.293 (34.1) 0.867 (1.27) 0.341 (37.2) 0.266 (31.4) 0.863 (1.42)
+ORPO 0.322 (29.4) 0.275 (26.2) 0.862 (0.69) 0.326 (31.4) 0.232 (14.6) 0.856 (0.64)
+DPO 0.285 (14.4) 0.242 (11.0) 0.869 (1.56) 0.278 (12.0) 0.208 (2.84) 0.867 (1.89)
+LC-DPO 0.283 (13.5) 0.259 (18.7) 0.871 (1.75) 0.299 (20.3) 0.248 (22.5) 0.871 (2.34)
+IPO BERTScore 0.283 (13.6) 0.247 (13.1) 0.870 (1.62) 0.282 (13.6) 0.218 (7.88) 0.866 (1.75)
+KTO 0.304 (21.9) 0.279 (28.2) 0.872 (1.88) 0.308 (24.0) 0.249 (23.1) 0.869 (2.11)
+ORPO 0.291 (16.9) 0.265 (21.4) 0.869 (1.57) 0.298 (19.9) 0.240 (18.4) 0.870 (2.21)
CheXagent-2 0.326 0.297 0.888 0.349 0.304 0.892
+DPO GREEN 0.387 (18.9) 0.339 (14.1) 0.891 (0.30) 0.387 (10.9) 0.320 (5.34) 0.888 (-0.38)
+DPO BERTScore 0.352 (8.11) 0.326 (9.58) 0.896 (0.95) 0.359 (2.88) 0.310 (1.78) 0.893 (0.15)
Table 3: Results on the MIMIC-CXR and CheXpert Plus test sets (percentage change compared to SFT baseline in brackets). Best results in bold, separate for CheXagent and CheXagent-2.
F1-14 F1-5
Model Judge Macro ()($\uparrow$)( ↑ ) Micro ()($\uparrow$)( ↑ ) Macro ()($\uparrow$)( ↑ ) Micro ()($\uparrow$)( ↑ ) Avg. ()($\uparrow$)( ↑ )
GPT-4V 20.4 35.5 19.6 25.8 25.3
MAIRA-1 38.6 55.7 47.7 56.0 49.5
MAIRA-2 41.6 58.1 50.4 59.1 52.3
Med-PaLM M (12B) 37.3 51.4 50.6 56.5 49.0
Med-PaLM M (84B) 39.8 53.6 51.6 57.9 50.7
Med-PaLM M (562B) 37.8 51.6 49.9 56.3 48.9
LLaVA-Rad 39.5 57.3 47.7 57.4 50.5
CheXagent 38.9 50.9 47.6 54.1 47.9
+DPO 41.5 (6.87) 54.1 (6.18) 51.8 (8.83) 58.3 (7.78) 51.4 (7.43)
+LC-DPO 37.6 (-3.32) 52.1 (2.20) 47.5 (-0.02) 55.2 (1.92) 48.1 (0.45)
+IPO GREEN 39.0 (0.39) 52.9 (3.87) 48.9 (2.88) 56.5 (4.36) 49.3 (3.06)
+KTO 40.7 (4.81) 55.0 (8.07) 51.6 (8.52) 59.3 (9.54) 51.7 (7.93)
+ORPO 41.4 (6.64) 55.0 (8.03) 51.5 (8.20) 58.3 (7.69) 51.5 (7.69)
+DPO 42.5 (9.37) 56.1 (10.1) 54.7 (15.0) 61.6 (13.8) 53.7 (12.2)
+LC-DPO 42.3 (8.75) 56.3 (10.6) 52.7 (10.8) 60.8 (12.4) 53.0 (10.8)
+IPO BERTScore 43.1 (10.8) 57.0 (11.9) 53.6 (12.6) 61.5 (13.6) 53.8 (12.4)
+KTO 44.0 (13.3) 58.0 (13.9) 54.0 (13.5) 62.3 (15.1) 54.6 (14.0)
+ORPO 42.4 (9.16) 56.7 (11.4) 52.5 (10.5) 60.7 (12.2) 53.1 (10.9)
CheXagent-2 44.6 57.8 55.5 62.4 55.1
+DPO GREEN 45.8 (2.6) 59.7 (3.3) 56.0 (0.8) 64.1 (2.7) 56.4 (2.4)
+DPO BERTScore 43.0 (-3.6) 59.7 (3.3) 53.5 (-3.6) 63.8 (2.3) 55.0 (-0.1)
Table 4: CheXbert scores on the MIMIC-CXR test set. Percentage change compared to SFT baseline in brackets. Best results in bold.
To further emphasize the issue observed for CheXagent with GREEN as Judge, we show the best configurations according to GREEN, and LC-GREEN, in Table 2. For CheXagent, using GREEN to guide the hyperparameter search leads to substantially more verbose reports for DPO, LC-DPO ($\alpha=0.001$), IPO, and KTO. Qualitative evaluation of the resulting reports indicates that the added verbosity is mostly due to exact, semantically or syntactically, repetitions. This is very likely a manifestation of reward hacking via length exploitation. To mitigate this issue, we use LC-GREEN instead of GREEN to guide the hyperparameter search. [^7] We opt for this approach, in lieu of obtaining new preference data and re-running the alignment due to computational constraints. As shown in Table 2, this leads to substantially less added verbosity. For CheXagent-2, it is less clear-cut whether length exploitation is present, as the highest and lowest scores are obtained with a similar verbosity. Note that this is in spite of a much more significant spread in average length of the reports in the chosen and rejected subsets. One possible explanation is that the SFT baseline produce substantially shorter reports than the references, and the increase in verbosity actually pushes the average closer to that of the references, effectively closing the gap rather than extending it by overshooting the average length of the references.
5.2 Judge Optimization Results
Results, using single runs, for GREEN, LC-GREEN, and the BERTScore on the MIMIC-CXR and CheXpert Plus test sets are available in Table 3. Additional results for ROUGE-L and BLEU-4 are available in Table 9. As expected, using GREEN as Judge results in the largest boost in GREEN, and using the BERTScore in the BERTScore. In particular, the top performing configuration on the MIMIC-CXR test set for CheXagent according to GREEN and LC-GREEN is KTO, using GREEN as Judge, boosting these metrics by 31.9% and 34.1% percent, respectively. For CheXagent, the top performing configuration according to the BERTScore is obtained by KTO, using the BERTScore as Judge. We observe similar trends for CheXagent-2 with DPO yielding an increase of 18.9% and 14.1% for GREEN and LC-GREEN, respectively when using GREEN as Judge. Notably, we observed overall similar trends for the CheXbert Plus, despite representing two different distributions: MIMIC-CXR was collected in a emergency department and CheXpert Plus was collected from in- and out-patient centers.
5.3 Generalization to CheXbert Scores
Accuracy ()($\uparrow$)( ↑ )
Model Judge View Classification Binary Image Single Disease Multi Disease Visual Question Image-Text Avg.
Classification Identification Identification Answering Reasoning
CheXagent 98.5 83.8 62.8 69.0 75.5 66.3 76.0
+DPO 98.5 84.0 63.2 68.7 75.5 66.1 76.0
+LC-DPO 98.5 83.7 62.2 68.3 75.5 65.3 75.6
+IPO GREEN 98.3 83.8 62.7 68.3 76.8 65.5 75.9
+KTO 98.3 83.8 63.2 68.7 74.6 65.0 75.6
+ORPO 98.2 83.5 63.5 68.6 74.3 64.2 75.4
+DPO 98.5 84.3 62.8 68.4 75.1 66.8 76.0
+LC-DPO 98.5 84.1 63.2 68.3 75.1 65.5 75.8
+IPO BERTScore 98.5 83.8 62.9 68.2 76.0 66.3 76.0
+KTO 98.5 84.2 63.2 68.5 76.3 66.3 76.2
+ORPO 98.3 84.1 62.7 68.8 73.9 66.3 75.7
CheXagent-2 99.0 83.0 65.5 83.9 83.2 78.7 82.2
+DPO GREEN 99.2 84.8 66.7 84.4 80.5 77.9 82.2
+DPO BERTScore 99.2 83.0 65.9 84.5 82.3 78.2 82.2
Table 5: Performance on six image perception and reasoning tasks different from RRG on the datasets listed in §4.2.
Although GREEN is a high quality and clinically relevant metric, the observed performance gains might be inflated due to the fact that we used GREEN as Judge for preference data generation. Thus, we instead turn to the CheXbert scores to be our silver-standard, a low-cost approximation of expert human (radiologists) judgment. CheXbert scores are ubiquitous in this setting and we include results from a representative sample of recent state-of-the-art medical VLMs for the RRG tasks: GPT-4V [^8] https://openai.com/index/gpt-4v-system-card/., MAIRA-1 (Hyland et al., 2023), MAIRA-2 (Bannur et al., 2024), Med-PaLM (Tu et al., 2024), and LLaVA-Rad (Chaves et al., 2024). We report macro and micro averaged F1-scores for the full 14 categories as well as the 5 categories subset. In addition, we provide the average across these scores. Results for our method are averages based on 1000 bootstrap samples, with confidence intervals available in Table 10.
For CheXagent, which is comparable to LLava-Rad and MAIRA-1 prior to alignment, our method boost the average CheXbert scores by up to 7-8% and 14% using GREEN and the BERTScore, respectively. Very interestingly, and unexpectedly, the top performing setup for CheXagent using the BERTScore as Judge is better than that for GREEN as Judge. One possible explanation for this is reward overoptimziation, as this seems to be a more prominent issue for GREEN than for the BERTScore. For CheXagent-2, which is the overall state-of-the-art prior to alignment and second only to MAIRA-2999Notably, MAIRA-2 use additional information at train and test time, including the radiologist-written prior report, when available. in terms of micro F1-14 scores, we can see that it is possible to improve performance even further using our proposed method. We note, however, that only the micro F1-14 scores are statistically significantly different from CheXbert-2, at a 10% level, for both GREEN and the BERTScore as Judge. Thus, we depict that it is possible to improve upon the already very strong micro averaged F1-scores, even when employing a general domain NLG metric like the BERTScore. Granular results, for each of the 14 categories, are available in Table 11.
5.4 Alignment Tax Analysis
While RLHF is powerful, it has been observed that it might lead to performance degradations or, forgetting (Askell et al., 2021; Ouyang et al., 2022)Ouyang et al. (2022) assessed such an alignment tax by evaluating the aligned policies on several natural language processing (NLP) benchmarks. Inspired by this, we benchmark the SFT baseline and the aligned policies on six diverse image perception and reasoning tasks using datasets listed in §4.2. Although there are some minor variations, on average, the performance matches that of the SFT baselines. Thus, our method substantially improves the quality of generated reports without compromising the quality of other image understanding tasks.
6 Conclusion
Our study highlights the significant potential of including preference fine-tuning in the post-training pipeline of medical VLMs. By developing an automated pipeline we effectively address the prohibitive cost of obtaining radiologist preferences at scale. Using our approach, we have shown that DAAs can substantially improve AI-generated radiology reports in clinically meaningful ways without additional radiologist feedback. Our approach achieves state-of-the-art performance on the MIMIC-CXR dataset while maintaining robust capabilities across diverse visual reasoning tasks. The surprising effectiveness of, inexpensive, general-domain NLG metrics for preference pair generation suggests a promising path forward for computationally efficient exploration of online alignment algorithms in this setting.
7 Limitations
Our work focuses on only two VLMs, CheXagent (Chen et al., 2024a) and CheXagent-2 (Chen et al., 2024b). Other VLMs, from different families and sizes, may behave differently. In addition, our study of CheXagent-2 was more limited in nature and should be considered more preliminary.
Moreover, we treat clinically relevant metrics such as GREEN and CheXbert scores as the silver standard. While these metrics are highly relevant, a thorough evaluation involving clinical experts, radiologists, should be conducted for this study to be considered complete. This is an important direction for future work.
In addition, our hyperparameter search is non-exhaustive and it is possible that the relative ranking of the methods considered would change with a more extensive search.
Finally, we restrict ourselves to only offline DAAs. This leaves out a range of very competitive alignment algorithms, including on-policy RL algorithms, as well as the online, or iterative, counterparts to the DAAs considered. The recent success of DeepSeek-R1 DeepSeek-AI et al. (2025) has led to a resurgence of interest in on-policy RL alignment algorithms. We believe that this is an very interesting avenue for future work.
Appendix A RLHF and DAAs
RLHF (Ziegler et al., 2020; Stiennon et al., 2020; Ouyang et al., 2022) is based on the constrained reward maximization objective
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}[R_{\psi}(x,y)]-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right],$ (1)
where $\mathbb{D}_{\text{KL}}$ is the Kullback-Leibler (KL) divergence and $\pi_{\text{ref}}$ is the reference policy. $R_{\psi}$ is the proxy reward model learned on a dataset of human preferences $\mathcal{D}=\{x^{(n)},y_{c}^{(n)},y_{r}^{(n)}\}_{n=1}^{N},$ where $y_{c}$ and $y_{r}$ denote the chosen and rejected completions for the prompt $x$, such that $y_{c}\succ y_{r}|x.$
Whilst extremely powerful, RLHF is computationally heavy, involves several steps, and can be tricky to implement in practice. Relatively recently, a new class of algorithms called DAAs (Rafailov et al., 2024) have become increasingly popular. [^10] In this paper, we use this terminology more loosely than in  Rafailov et al. (2024). This class of algorithms re-parameterize the reward model via a change-of-variables using the closed-formed solution to the objective in (1), effectively bypassing both the reward modeling and reinforcement learning (RL) stages. Resulting in algorithms that remain performant yet computationally more light weight and easier to implement. DPO (Rafailov et al., 2023) was the first in this category and remains one of the most popular versions.
DPO exploits the closed-formed solution to (1), $\pi(y|x)\propto\pi_{\text{ref}}(y|x)\exp(R(x,y)/\beta)$ and the Bradley-Terry (BT) model (Bradley and Terry, 1952) of human preferences $p^{*}(y_{1}\succ y_{2}|x)=\sigma(\exp(R^{*}(x,y_{1}))-\exp(R^{*}(x,y_{2}))),$ where $R^{*}$ is the latent reward model, $\exp$ is the exponential function, and $\sigma$ is the logistic function. The reward can be isolated and written as a function of the policy $R(x,y)=\beta\log\frac{\pi(y|x)}{\pi_{\text{ref}}(y|x)}.$ This re-parametrization can be applied to the latent reward $R^{*}$ and substituted into the BT model, $p^{*}(y_{1}\succ y_{1}|x)=\sigma\left(\beta\log\frac{\pi^{*}(y_{1}|x)}{\pi_{\text{ref}}(y_{1}|x)}-\beta\log\frac{\pi^{*}(y_{2}|x)}{\pi_{\text{ref}}(y_{2}|x)}\right),$ where $\pi^{*}$ is the optimal policy corresponding to the latent reward. Crucially, the probability of human preferences is now in terms of the policy instead of the reward model. A parameterized policy $\pi_{\theta}$ can then be learned via a simple classification loss over the preference data
$\mathcal{L}_{\text{DPO}}(\theta)=-\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{c}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{c}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{r}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{r}\mid x\right)}\right).$
Hence, this change-of-variables has transformed a loss over rewards into a loss over policies.
Appendix B Implementation Details
B.1 Data Details
We use the official train, validation, and test splits for the MIMIC-CXR (Johnson et al., 2019) and CheXpert PlusChambon et al. (2024) datasets. A key difference between CheXagent and CheXagent-2 is that CheXagent-2 additionally use indications, which provide clinical context, to aid in the RRG task. This means that CheXagent-2 imposes a stricter data requirement (i.e. each image-report pair must have a corresponding indication). Due to this, CheXagent-2 has slightly fewer examples than CheXagent for the RRG task. The number of examples in each split is available in Table 6. For MIMIC-CXR, the difference in test and validation is so small that it is negligible for the RRG task. On the other hand, for CheXpert Plus the difference is quite substantial. However, this is immaterial for any conclusions in this paper as we do not directly compare the models with each other on the CheXpert Plus data.
The view classification, binary image classification, single disease identification, multi disease identification, VQA, and image-text reasoning test data from the RSNA (Shih et al., 2019), SIIM (American College of Radiology, 2019), OpenI (Demner-Fushman et al., 2016), SLAKE (Liu et al., 2021), and Rad-Restruct (Pellegrini et al., 2023) datasets are processed as in  Chen et al. (2024b).
CheXagent CheXagent-2
Dataset Train Validation Test Train Validation Test
MIMIC-CXR 148463 1164 2309 148090 1162 2300
CheXpert Plus 208 151
Table 6: Number of examples in each split of MIMIC-CXR and CheXpert plus for CheXagent and CheXagnet-2.
B.2 Training Details
All models are trained using a machine with 4xA100, 4xA6000 or 8xH100 GPUs using learning rate learning rate $10^{-6}$. We set global batch size to 32 for CheXagent (Chen et al., 2024a) and 64 for CheXagent-2 (Chen et al., 2024b). Each model is trained for one epoch. The image encoder is frozen while we train the LLM. Due to compute constraints, we only tune hyperparameters that are specific of the DAAs considered while keeping everything else fixed. An overview is given in Table 7, including optimal configurations. This non-exhaustive search was based on previous work and our initial experiments. For CheXagent, each $\lambda\in[0.5,1.0,4.0,5.0]$, ORPO resulted in a model which produced a special token at odd places, leading to a crash of our evaluation pipeline. We address this by catching the error and set the special token to the padding token. The same issue emerged for some $\lambda$ for CheXagent-2.
Algorithm Objective Hyperparameters
DPO (Rafailov et al., 2023) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{c}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{c}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{r}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{r}\mid x\right)}\right)$ $\beta\in[0.01,0.05^{*,\dagger},0.1]$
LC-DPO (Park et al., 2024) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{c}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{c}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{r}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{r}\mid x\right)}+\alpha(|y_{c}|-|y_{r}|)\right)$ $\beta\in[0.01,0.05^{*,\dagger},0.1],\alpha\in[0.001,0.01^{*,\dagger}]$
IPO (Azar et al., 2023) $\left(\log\frac{\pi_{\theta}(y_{c}|x)}{\pi_{\text{ref}}(y_{c}|x)}-\log\frac{\pi_{\theta}(y_{r}|x)}{\pi_{\text{ref}}(y_{r}|x)}-\frac{1}{2\tau}\right)^{2}$ $\tau\in[0.1,0.5,1.0^{*,\dagger}]$
KTO (Ethayarajh et al., 2024) $-\lambda_{c}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{c}|x)}{\pi_{\text{ref}}(y_{c}|x)}-z_{\text{ref}}\right)+\lambda_{r}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{r}|x)}{\pi_{\text{ref}}(y_{r}|x)}\right)$ $\beta\in[0.01,0.05^{*},0.1^{\dagger}],\lambda_{c}=\lambda_{r}$
$\text{where}\,\,z_{\text{ref}}=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\beta\mathbb{D}_{\text{KL}}\left(\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right)]$
ORPO (Hong et al., 2024) $-\log p_{\theta}(y_{c}|x)-\lambda\log\sigma\left(\log\frac{p_{\theta}(y_{c}|x)}{1-p_{\theta}(y_{c}|x)}-\log\frac{p_{\theta}(y_{r}|x)}{1-p_{\theta}(y_{r}|x)}\right),$ $\lambda\in[0.5,1.0,4.0^{\dagger},5.0^{*}]$
$\text{where}\,\,p_{\theta}(y|x)=\exp\left(\frac{1}{|y|}\log\pi_{\theta}(y|x)\right)$
Table 7: Hyperparameter search for all direct alignment algorithms (DAAs) considered in this paper. For CheXagent, and denotes best for GREEN and the BERTScore as Judge, respectively. For CheXagent-2, $\beta=0.1$ was found optimal for both GREEN and the BERTScore.
B.3 Sampling Details
In this paper, we set up CheXagent to treat the cases where the two CXRs are a frontal and a lateral image (i.e. from the same point in time) and a frontal and a prior frontal image (i.e. from two points in time) as separate cases with a separate prompt. This is in contrast to CheXagent-2, which employs the same prompt for both cases. Moreover, we employed stochastic sampling with temperature $1.0$ for CheXagent at test time, while CheXagent-2 is based on greedy sampling, both with beam size set to $1$. For preference pairs generation, both employ stochastic sampling with temperature $1.0$ and beam size set to $1$ These differences will explain parts of the performance differences between CheXagent and CheXagent-2. However, this is of minor concern to this study as we are mainly interested in the effect of preference fine-tuning, comparing the same setup prior and post alignment.
B.4 Evaluation Details
GREEN (Ostmeier et al., 2024) and GREEN-LC are based on the official implementation [^11] https://github.com/Stanford-AIMI/GREEN using StanfordAIMI/GREEN-radllama2-7b. The BERTScore (Zhang et al., 2019) used is from evaluate (v0.4.0) with distilbert-base-uncased as BERT model. The F1 CheXbert (Smit et al., 2020) scores use f1chexbert (v0.0.2) with default configurations. ROUGE (Lin, 2004) is from rouge-score (v0.1.2) using rougeL. Finally, the BLEU (Papineni et al., 2002) is BLEU-4, based on a custom code, available in the code base corresponding to this project.
CheXbert scores for GPT-4V [^12] https://openai.com/index/gpt-4v-system-card/., MAIRA-1 (Hyland et al., 2023), Med-PaLM (Tu et al., 2024), and LLaVA-Rad (Chaves et al., 2024) are borrowed directly from Supplementary Table 1 in Chaves et al. (2024). The CheXbert scores for MAIRA-2 (Bannur et al., 2024) are from Table D.1 in Bannur et al. (2024).
Appendix C Additional results
C.1 Summary Statistics of Preference Data
GREEN BERTScore
Metric Report Length Metric Report Length
CheXagent Mean Median Std. Mean Median Std. Mean Median Std. Mean Median Std.
Chosen 0.63 0.60 0.24 56.3 54.0 20.2 0.90 0.90 0.03 55.1 53.0 20.2
Rejected 0.26 0.22 0.19 52.7 51.0 24.9 0.85 0.86 0.04 52.1 50.0 25.3
CheXagent-2 Mean Median Std. Mean Median Std. Mean Median Std. Mean Median Std.
Chosen 0.57 0.50 0.25 0.24 0.20 0.20 0.91 0.90 0.03 0.85 0.86 0.05
Rejected 54.4 52.0 19.3 45.6 44.0 22.8 52.7 51.0 18.6 45.0 43.0 23.8
Table 8: Summary statistics of reference-based metric and report length in the chosen and rejected subsets using GREEN and the BERTScore as Judges.
C.2 Judge Optimization Results
MIMIC-CXR
Model Judge GREEN ()($\uparrow$)( ↑ ) GREEN-LC ()($\uparrow$)( ↑ ) BERTScore()($\uparrow$)( ↑ ) BLEU-4 ()($\uparrow$)( ↑ ) ROUGE-L ()($\uparrow$)( ↑ )
CheXagent 0.249 0.218 0.856 0.041 0.274
+DPO 0.325 (30.6) 0.263 (20.5) 0.862 (0.65) 0.057 (39.9) 0.293 (6.94)
+LC-DPO 0.320 (28.7) 0.288 (32.0) 0.864 (0.90) 0.056 (38.4) 0.302 (10.4)
+IPO GREEN 0.326 (30.9) 0.282 (29.3) 0.863 (0.84) 0.059 (45.0) 0.297 (8.57)
+KTO 0.328 (31.9) 0.293 (34.1) 0.867 (1.27) 0.056 (38.8) 0.305 (11.4)
+ORPO 0.322 (29.4) 0.275 (26.2) 0.862 (0.69) 0.053 (29.5) 0.290 (6.08)
+DPO 0.285 (14.4) 0.242 (11.0) 0.869 (1.56) 0.059 (44.2) 0.309 (12.9)
+LC-DPO 0.283 (13.5) 0.259 (18.7) 0.871 (1.75) 0.057 (41.2) 0.315 (15.1)
+IPO BERTScore 0.283 (13.6) 0.247 (13.1) 0.870 (1.62) 0.060 (47.2) 0.309 (12.9)
+KTO 0.304 (21.9) 0.279 (28.2) 0.872 (1.88) 0.056 (37.7) 0.310 (13.4)
+ORPO 0.291 (16.9) 0.265 (21.4) 0.869 (1.57) 0.054 (33.8) 0.307 (12.0)
CheXagent-2 0.326 0.297 0.888 0.136 0.350
+DPO GREEN 0.387 (18.9) 0.339 (14.1) 0.891 (0.301) 0.150 (10.3) 0.357 (1.93)
+DPO BERTScore 0.352 (8.11) 0.326 (9.58) 0.896 (0.949) 0.153 (12.5) 0.372 (6.37)
CheXpert Plus
Model Judge GREEN ()($\uparrow$)( ↑ ) GREEN-LC ()($\uparrow$)( ↑ ) BERTScore()($\uparrow$)( ↑ ) BLEU-4 ()($\uparrow$)( ↑ ) ROUGE-L ()($\uparrow$)( ↑ )
CheXagent 0.248 0.202 0.851 0.038 0.274
+DPO 0.322 (29.9) 0.222 (9.69) 0.859 (0.91) 0.053 (40.2) 0.289 (5.25)
+LC-DPO 0.330 (32.8) 0.268 (32.3) 0.861 (1.26) 0.049 (30.2) 0.305 (10.6)
+IPO GREEN 0.334 (34.6) 0.267 (31.9) 0.861 (1.15) 0.052 (39.1) 0.295 (7.22)
+KTO 0.341 (37.2) 0.266 (31.4) 0.863 (1.42) 0.052 (36.9) 0.305 (10.9)
+ORPO 0.326 (31.4) 0.232 (14.6) 0.856 (0.64) 0.046 (21.9) 0.287 (4.63)
+DPO 0.278 (12.0) 0.208 (2.84) 0.867 (1.89) 0.054 (43.6) 0.307 (11.3)
+LC-DPO 0.299 (20.3) 0.248 (22.5) 0.871 (2.34) 0.056 (47.6) 0.318 (15.1)
+IPO BERTScore 0.282 (13.6) 0.218 (7.88) 0.866 (1.75) 0.052 (37.6) 0.309 (12.0)
+KTO 0.308 (24.0) 0.249 (23.1) 0.869 (2.11) 0.054 (43.8) 0.316 (14.6)
+ORPO 0.298 (19.9) 0.240 (18.4) 0.870 (2.21) 0.049 (30.8) 0.314 (14.0)
CheXagent-2-3b 0.349 0.304 0.892 0.123 0.361
+DPO GREEN 0.387 (10.89) 0.320 (5.34) 0.888 (-0.37) 0.126 (2.74) 0.356 (-1.30)
+DPO BERTScore 0.359 (2.88) 0.310 (1.78) 0.893 (0.15) 0.130 (2.86) 0.357 (-0.91)
Table 9: Results on the MIMIC-CXR and CheXpert Plus test sets (percentage change compared to SFT baseline in brackets). Best results in bold, separate for CheXagent and CheXagent-2.
C.3 Generalization to CheXbert Scores
F1-14 F1-5
Model Judge Macro ()($\uparrow$)( ↑ ) Micro ()($\uparrow$)( ↑ ) Macro()($\uparrow$)( ↑ ) Micro ()($\uparrow$)( ↑ ) Avg. ()($\uparrow$)( ↑ )
GPT-4V 20.4 35.5 19.6 25.8 25.3
MAIRA-1 38.6 55.7 47.7 56.0 49.5
MAIRA-2 41.6 58.1 50.4 59.1 52.3
Med-PaLM M (12B) 37.3 51.4 50.6 56.5 49.0
Med-PaLM M(84B) 39.8 53.6 51.6 57.9 50.7
Med-PaLM M (562B) 37.8 51.6 49.9 56.3 48.9
LLaVA-Rad 39.5 57.3 47.7 57.4 50.5
CheXagent 38.9[37.8,40.0] 50.9[50.0,51.8] 47.6[46.3,48.8] 54.1[52.9,55.3] 47.9[46.9,48.8]
[…]
Table 10: CheXbert scores on the MIMIC-CXR test set. 90% confidence interval obtained by 1000 bootstrap samples in subscripts. Best results in bold.
F1-scores ()($\uparrow$)( ↑ )
Model Judge ECm. Cmgl. LOpac. LLes. Edema Cnsl. Pna. Atel. PEff. Pmtx. POth. Frac. SuDev. NoF.
CheXagent 0.347 0.620 0.461 0.171 0.493 0.158 0.227 0.453 0.655 0.444 0.092 0.240 0.787 0.304
+DPO 0.372 0.675 0.444 0.186 0.525 0.226 0.231 0.452 0.710 0.416 0.178 0.217 0.822 0.365
+LC-DPO 0.385 0.666 0.380 0.178 0.430 0.211 0.158 0.404 0.665 0.407 0.083 0.129 0.835 0.339
+IPO GREEN 0.402 0.686 0.395 0.160 0.472 0.222 0.206 0.405 0.661 0.483 0.090 0.138 0.822 0.321
+KTO 0.388 0.682 0.422 0.202 0.542 0.192 0.180 0.444 0.721 0.552 0.074 0.118 0.841 0.349
+ORPO 0.348 0.684 0.479 0.201 0.492 0.224 0.247 0.475 0.698 0.511 0.072 0.177 0.835 0.365
+DPO 0.373 0.688 0.467 0.192 0.580 0.236 0.141 0.503 0.728 0.483 0.157 0.239 0.831 0.336
+LC-DPO 0.301 0.692 0.506 0.208 0.567 0.175 0.107 0.471 0.731 0.567 0.167 0.225 0.832 0.376
+IPO BERTScore 0.386 0.690 0.500 0.218 0.559 0.190 0.111 0.500 0.739 0.550 0.195 0.220 0.836 0.342
+KTO 0.430 0.701 0.487 0.217 0.577 0.172 0.201 0.514 0.735 0.584 0.166 0.197 0.830 0.360
+ORPO 0.377 0.690 0.491 0.148 0.562 0.157 0.231 0.481 0.736 0.497 0.116 0.218 0.826 0.418
CheXagent-2 0.324 0.672 0.539 0.244 0.606 0.235 0.284 0.543 0.720 0.527 0.085 0.330 0.815 0.333
+DPO 0.450 0.696 0.493 0.218 0.606 0.214 0.325 0.545 0.738 0.526 0.157 0.255 0.856 0.343
+DPO 0.358 0.703 0.547 0.166 0.622 0.127 0.066 0.478 0.749 0.550 0.143 0.271 0.853 0.396
Table 11: F1 scores on the MIMIC-CXR test set using 14 categories from the CheXbert labeler (Smit et al., 2020): Enlarged Cardiomediastinum (ECm.), Cardiomegaly (Cmgl.), Lung Opacity (LOpac.), Lung Lesion (LLes.), Edema, Consolidation (Cnsl.), Pneumonia (Pna.), Atelectasis (Atel.), Pleural Effusion (PEff.), Pneumothorax (Pmtx.), Pleural Other (POth.), Fracture (Frac.), Support Devices (SuDev.), no Findings (NoF.). Best results in bold, separate for CheXagent and CheXagent-2.